OpenStudy (anonymous):

i chose C am i right? 2. A store allows customers to fill their own bags of candy. Glen decides he only wants caramels and candy corn. Caramels sell for $0.95 per pound, and candy corn sells for $0.79 per pound. Glen’s bag weighs 1.7 pounds and it costs $1.48. If p = pounds of caramels, which equation in simplest form represents the situation described. (Points : 1) –0.16p + 1.62 = 1.48 –0.16p + 1.62= 1.7 0.16p + 1.34 = 1.48 0.16p + 1.34 = 1.7

OpenStudy (anonymous):

3. A store allows customers to fill their own bags of candy. Glen decides he only wants caramels and candy corn. Caramels sell for $0.95 per pound, and candy corn sells for $0.79 per pound. Glen’s bag weighs 1.7 pounds and it costs $1.48. How many pounds of caramels and candy corn did Glen buy? (Points : 1) 0.8 pounds of caramels; 0.9 pounds of candy corn 0.9 pounds of caramels; 0.8 pounds of candy corn 1 pound of caramels; 0.7 pounds of candy corn 0.7 pounds of caramels; 1 pound of candy corn 4. Melania has two jobs. During the day, she works as an office clerk, and in the evening she works as a cashier. Her office job pays her $13 per hour, and her cashier job pays her $9.25 per hour. In one week, Melania worked 38 hours, and she earned a total of $434. If c represents the number of hours that Melania works as an office clerk, write an equation that can be used to find how many hours Melania worked in both jobs. (Points : 1) 13c + 9.25(c – 38) = 434 13c + 9.25(38 – c) = 434 13c + 9.25(434 – c) = 38 9.25c + 13(38 – c) = 434 5. Soren has two jobs. During the day, he works as an office clerk, and in the evening he works as a cashier. His office job pays him $11 per hour, and his cashier job pays him $8.75 per hour. In one week, Soren worked 46 hours, and he earned a total of $470. Find the number of hours Soren worked in each job. (Points : 1) Office clerk: 16 hours; cashier: 30 hours Office clerk: 20 hours; cashier: 26 hours Office clerk: 30 hours; cashier: 16 hours Office clerk: 31 hours; cashier: 15 hours

OpenStudy (anonymous):

3. C 4. C 5. D

OpenStudy (anonymous):

Yes you are correct! Good Job!!

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

No Problem :-)

OpenStudy (anonymous):

I got 3 wrong :(